solve for x : x 7 = 1 { the answer is not only = 1, it probably has 6 other values }

solve for x :


x7 = 1


 


{ the answer is not only = 1, it probably has 6 other values }

Grade:12

2 Answers

Askiitian.Expert Rajat
14 years ago

Hi Anthony

1 = cos2pi + isin2pi

11/7 = cos(2rpi/7) + isin(2rpi/7)

where

r = 1,2,3,4,5,6,7

This was derived using Euler's Formula

If you want you can join Askiitians.com's Online IITJEE Coaching Sessions.

For more details, visit the Packages Section of Askiitians.com

 

Regards,

Rajat

Askiitians.com

anthony rebello
14 years ago

thnx for the reply, but i hav a doubt :

 

why cant    1=cos (0) + i sin (0)